Test the following series for convergence


2.3/(4^2.5^2)+ 4.5/6^2.7^2) + 6.7/ (8^2.9^2) ..+.......

1
Expert's answer
2021-11-14T18:11:31-0500

Let us test the series "\sum\limits_{n=1}^{\infty}\frac{2n(2n+1)}{(2n+2)^2(2n+3)^2}" for convergence. Let us consider the series "\sum\limits_{n=1}^{\infty}\frac{1}{n^2}."

Since "\lim\limits_{n\to\infty}\frac{\frac{2n(2n+1)}{(2n+2)^2(2n+3)^2}}{\frac{1}{n^2}}
=\lim\limits_{n\to\infty}\frac{2n^3(2n+1)}{(2n+2)^2(2n+3)^2}
=\lim\limits_{n\to\infty}\frac{4+\frac{2}{n}}{(2+\frac{2}n)^2(2+\frac{3}n)^2}
=\frac{4}{2^22^2}=\frac{1}4,"

we conclude that the series "\sum\limits_{n=1}^{\infty}\frac{2n(2n+1)}{(2n+2)^2(2n+3)^2}" is equivalent to the series "\sum\limits_{n=1}^{\infty}\frac{1}{n^2}". Since the "s"-series "\sum\limits_{n=1}^{\infty}\frac{1}{n^s}" is covergent for "s>1," we conclude that the series "\sum\limits_{n=1}^{\infty}\frac{2n(2n+1)}{(2n+2)^2(2n+3)^2}" is covergent as well.